OpenStudy (anonymous):

I'm a bit confused on how to do this I will medal and fan if someone can answer and explain this. Using the points (3, 4) and (-5, 8) find the following: a. Slope. b. Point-Slope form of the equation. c. Slope-intercept form of the equation.

OpenStudy (anonymous):

Slope formula is \[m=\frac{ y2-y1 }{ x2-x1 }\] So you plug in (8-4)/(-5-3) and that's your slope for A. :)

OpenStudy (anonymous):

For B, point slope form is y-y1=m(x-x1) So basically, you plug in y-4=m(x-3) With your slope in there too.

OpenStudy (anonymous):

For C, slope-intercept form is y=mx+b So you can take your point-slope y-4=m(x-3) <--with slope in there too The slope needs to distribute to what's in the parentheses, so y-4=mx-3m Then y needs to be by itself, so add the 4 to both sides. y=mx-3m+4 Hope that all helps! If you're still confused I can work it out for you, or just check your answers :)

OpenStudy (anonymous):

@bruhhh Thank you!